OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x - y = 3 7x - 9y = -2

OpenStudy (anonymous):

multiply the first by -9 -36x+9y=-27 7x-9y=-2 add em -29x=-29 x=1 7(1)-9y=-2 -9y=-9 y=1 Answer's (1,1)

OpenStudy (saifoo.khan):

1,1 is right answer.

OpenStudy (anonymous):

The first equation gives us\[4x-y=3\]\[y = 4x -3\] Substitute into the second equation: \[7x - 9(4x-3) = -2\]\[7x-36x+27=-2\]\[-29x=-29\]\[x=1\]\[y=1\] (the last step from equation 1)

OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + y = 3 2x + 2y = 10 (2 points)Question 5 options: 1) (1, 2) 2) (2, 3) 3) No Solutions 4) Infinitely Many Solutions

OpenStudy (anonymous):

No solutions: \[2x + 2y =2(x+y) = 2(3) = 6\] So the second equation is wrong. (Step 3 done by substituting in equation 1)